To solve a circuit means to use the values of potentials and resistances to find all currents passing through, and voltage drops across all elements in the circuit. Create a general procedure to solve any circuit DC circuit with any number of resistors.

Consider the circuit shown below. Find I1, V1, I2, and V3.

1 Aww R1-1.0012 6.001 13.00 Ω
Answer #1

R2 and R3 are in parallel and their combination is given as

R23 = R2 R3/(R2 + R3) = 6 x 13/(6 + 13) = 4.11 ohm

R1 and R23 are in series and their combination is given as

R123 = R1 + R23 = 1 + 4.11 = 5.11 ohm

V = battery Voltage = 12 Volts

the battery is connected in series with resistor R1 , hence the current in resistor R1 is same as the total current coming from the battery and hence is given using ohm's law as

i1 = itotal = V/R123

i1 = 12/5.11

i1 = 2.35 A

V1 = potential difference across R1 = i1 R1 = (2.35) (1) = 2.35 Volts

The potential difference across parallel resistors remains equal . hence

V2 = V3 = V - V1

V2 = V3 = 12 - 2.35 = 9.65 volts

i2 = current in R2 = V2/R2 = 9.65/6 = 1.6 A

i3 = current in R3 = V3/R3 = 9.65/13 = 0.74 A